A consumer’s utility function is given by π‘ˆ =∝ 𝐼𝑛π‘₯1 + 𝛽𝐼𝑛π‘₯2

. Find the values of π‘₯1

andΒ 

π‘₯1 which maximize utility subject to the budgetary constraint 𝑃1π‘₯1 + 𝑃2π‘₯2 = 𝑀


1
Expert's answer
2021-02-25T18:16:30-0500
"\frac {\delta U}{\delta x_1}=\frac {\alpha}{x_1}"


"\frac {\delta U}{\delta x_2}=\frac {\beta}{x_2}"


"\frac {\alpha}{x_1p_1}=\frac {\beta}{x_2p_2}"

"x_2p_2=\frac{\beta}{\alpha}x_1p_1"

"x_1=\frac {2M}{p_1(\alpha+\beta)}"